OpenStudy (anonymous):

If u and v are the vectors below, find the vector w whose tail is at the point halfway from the tip of v to the tip of u and whose head is at the point halfway from the tip of u to the tip of u-v.

OpenStudy (anonymous):

I can see the vectors u and v plz

OpenStudy (anonymous):

u = -1 3 4 v = 2 4 -3

OpenStudy (anonymous):

and provide an explanation please and thank you!

OpenStudy (anonymous):

x coordinate of vector w is: (-1+2)/2=0.5; y coordinate of vector w is: (3+4)/2=3.5; z coordinate of vector w is: (4-3)/2=0.5; I think these coordinates make up the vector w but its probably wrong.

OpenStudy (anonymous):

basically I just averaged the x y and z coordinates of u and v to get w. I think you have to exchange the z coordinates with the x coordinates but I'm not sure. Good luck!